How long it takes for 600 to amount 900 at an annual rate of 6% compound quarterly


1
Expert's answer
2021-11-19T07:08:37-0500

Investment = 600

Future Value = 900

Quarterly Compounding

Rate = 6%/4=1.5%perquarter


"FutureValue=Investment×(1+InterestRate)^N\\

FutureValue=Investment×(1+InterestRate)^N \\

900=600×(1.015)^N \\"

"900=600×(1.015)^N \\

1.5=(1.015)^N"

Taking log on both sides

Log(1.5)/Log(1.015)=N

N = 27.23 quarters or 6.80 years